Find the slop and y-intercept for each line.
y = -x + 5
3x + 5y + 10 = 0

Answer by Alan3354(69443)

Find the slop and y-intercept for each line.
y = -x + 5
slope = -1
Y-int = 5
---------------
3x + 5y + 10 = 0
slope = -3/5
y-int = -2